# Human or LLM as Standardized Patients? A Comparative Study for Medical Education

### v1

@@ Abstract
Standardized Patients (SP) are indispensable for clinical skills training but remain expensive, inflexible, and difficult to scale. Existing large-language-model (LLM)–based SP simulators promise lower cost yet show inconsistent behavior and lack rigorous comparison with human SP. We present EasyMED, a multi-agent framework combining a Patient Agent for realistic dialogue, an Auxiliary Agent for factual consistency, and an Evaluation Agent that delivers actionable feedback. To support systematic assessment, we introduce SPBench, a benchmark of real SP–doctor interactions spanning 14 specialties and eight expert-defined evaluation criteria. Experiments demonstrate that EasyMED matches human SP learning outcomes while producing greater skill gains for lower-baseline students and offering improved flexibility, psychological safety, and cost efficiency.
@@
A Comparative Study for Medical Education
@@ 1 Introduction
Medical education underpins professional healthcare training, fostering essential competencies in clinical reasoning, communication, and decision-making cleland2022researching. Deficiencies in these skills contribute to diagnostic errors and preventable harm, with communication failures alone responsible for thousands of in-hospital deaths annually ball2015improving. To improve learning efficiency and scalability, recent studies have explored artificial intelligence (AI)—particularly LLM—for clinical education. LLM enable interactive simulation and automated feedback at low cost, offering new possibilities for scalable, practice-oriented training  thirunavukarasu2023large.
Patient simulation is central to such experiential learning, allowing students to practice reasoning and communication in realistic but safe settings. The Standardized Patient—actors trained to portray clinical cases—is widely recognized as the gold standard  sayers2024database; ma2023standardized, yet remains costly, labor-intensive, and difficult to scale, limiting case diversity and accessibility zendejas2013cost.
To mitigate these constraints, previous researchers have explored LLM-based virtual SP simulators as low-cost and scalable alternatives ye2025multimodal; du2024llms; khandwaha2025designing. Most systems adopt a single-agent, prompt-driven architecture, where a pretrained model performs patient simulation, intent understanding, and dialogue control within one process. This unified design couples persona control, information disclosure, and evaluation, leading to unstable role behavior, factual errors, and uncontrolled information flow sirdeshmukh2025multichallenge; laban2025llms. Evaluation methods are usually ad hoc, based on coarse rubrics or single-turn scoring, which limits analysis of interactive reasoning and learning outcomes burke2024assessing; wang2025survey. As a result, whether virtual SP can match the performance and educational value of human SP remains unclear simzine2025vsp.
In this research, we introduce EasyMED, a multi-agent framework that transforms virtual SP training into a structured and interactive process. Unlike traditional single-agent models li2023chatdoctor, EasyMED decomposes patient simulation, intent recognition, and evaluation into coordinated agents with distinct roles. Each agent aligns with the workflow of real clinical practice: the Patient Agent simulates realistic symptom expression and emotional behavior, the Auxiliary Agent interprets the doctor’s conversational intent to guide contextually appropriate responses, and the Evaluation Agent monitors and assesses interactions to ensure consistency and clinical validity. This design establishes an interpretable framework that enhances controllability and error traceability through clearly defined, task-specific agent roles.
To benchmark interactive behaviors that matter pedagogically, we introduce SPBench, a dataset derived from authentic SP-student dialogues covering 14 specialties and eight expert-defined assessment criteria. It standardizes prompts, information disclosure rules, and scoring procedures, enabling consistent comparison of models and prompting strategies under realistic clinical scenarios.
We further validate EasyMED through a four-week controlled study comparing it with human SP training under identical content and scoring rubrics. The results demonstrate that EasyMED achieves comparable or even superior learning outcomes—particularly among novice learners—while providing greater flexibility, psychological safety, and cost efficiency.
In summary, our contributions are threefold: (1) Multi-Agent Framework for Virtual SP Training: We propose EasyMED, effectively simulates the structured interactive process of the realistic SP and training doctors with three coordinated agents. This structure enables controllable, persona-stable interactions and provides transparent, multi-dimensional feedback. (2) Benchmarking with SPBench: We construct a benchmark from authentic SP–student dialogues covering 14 specialties and eight evaluation criteria. It standardizes prompts, disclosure rules, and scoring, supporting reproducible, turn- and session-level evaluation of interactive behaviors. (3) Empirical Validation of Educational Impact: We conduct a four-week controlled study comparing EasyMED with human SP training under identical test content and rubrics, demonstrating comparable learning outcomes, improved accessibility, and clear advantages for lower-baseline learners.
2 Background and Motivation
2.1 The Definition of Standardized Patients
Definition
Standardized Patients (SP) is a training approach in which individuals are trained to consistently simulate the history, personality, physical findings, and emotional states of real patients. SP is widely used in medical education, especially in Objective Structured Clinical Examinations (OSCE), providing a safe and repeatable environment for practicing key clinical skills such as history taking, physical examination, and communication cook2009virtual; berman2016role; kononowicz2015virtual. This allows students to develop these abilities without risking harm to real patients.
Challenge
While SP is effective, it faces several challenges, including high costs for recruiting and training actors, logistical constraints from fixed locations and schedules, and inconsistent performance due to factors like fatigue or mood hamilton2024evolution; kuehne2018standardized. These limitations restrict the scalability and accessibility of SP-based training, particularly in resource-limited settings.
2.2 LLM Simulator vs. Standardized Patient
As discussed in Section 2.1, the high cost, limited scalability, and inconsistent performance of traditional SP have prompted researchers to explore LLM-based virtual SP as an alternative. While LLM-based SP reduce costs, improve scalability, and standardize responses compared with human SP, they still show unstable behavior, lack structured feedback, and have limited validation against human SP. Building on these questions, our EasyMED framework further enhances virtual SP capabilities by introducing a multi-agent design, structured case control, and integrated evaluation modules.
Metric Human SP Existing LLM-based SP Our EasyMED
Cost High actor, training and venue costs Low API costs; limited assessment Low costs with built-in multi-dimensional assessment
Time Fixed scheduling slots Instant but single-agent Instant with multi-agent coordination
System Architecture Single-agent role-play; no module separation wang2023notechat; wang2024patient Multi-agent with Patient, Assistant, Evaluation
Assessment Capability Manual checklists Manual or simple scoring; no integrated module li2024leveraging Built-in Evaluation Agent with automated, multi-dimensional feedback
Benchmarking No real-case benchmark; mostly synthetic sardesai2024utilizing; ye2025multimodal SPBench: real SP–student dialogues (12 specialties, 8 dimensions)
Empirical Validation Local experience only Few or no long-term comparative studies ye2025multimodal Four-week study comparing EasyMED vs. Human SP
Persona Stability Actor variability Persona inconsistency, leakage, harmful replies du2024llms Prompt control + multi-agent improves stability
Consistency Performance varies across actors Standardized but unstable persona Multi-agent design improves persona stability
Feedback Manual scoring Limited or no automated feedback Evaluation Agent + SPBench for reliable feedback
Reproducibility Limited, hard to repeat Anytime but limited diversity Anytime with curated real-case benchmark
Psychological Safety Higher performance anxiety Lower pressure vs. Human SP almansoori2025self. Lower pressure + structured, transparent assessment
Table 1: Integrated comparison of Human SP, existing LLM-based SP systems, and our EasyMED framework across operational efficiency, system design, teaching quality, and learner experience.
Figure 1: The multi-agent architecture of our virtual medical training system. This “division of labor” model includes: (1) a Patient Agent to simulate patient dialogue, (2) an Intention Identification Agent to interpret the user’s clinical queries, and (3) an Evaluation Agent to assess performance against ground truth.
As shown in Table 1, these limitations are evident when comparing Human SP, existing LLM-based SP systems, and our EasyMED framework across operational efficiency, system design and validation, teaching quality, and learner experience.
3 Methodology
3.1 EasyMED: Multi-Agent Framework
Building on the limitations of single-agent simulators discussed in Section 1, EasyMED is designed as a controllable and modular training framework. Existing virtual SP systems integrate patient simulation, intent understanding, and evaluation within one model, which often leads to persona drift, uncontrolled information disclosure, and weak or non-actionable feedback du2024llms. To address these issues, EasyMED adopts a multi-agent division-of-labor architecture (Figure 1) that explicitly separates patient simulation, intent interpretation, and evaluation, while maintaining coordination within a single session.
During a training session, a student’s utterance is first analyzed by an Auxiliary Agent, which identifies the clinical intent behind the question and maps it to one of 31 expert-defined intent categories (see Appendix 9 for qualifications). This output conditions the Patient Agent, which generates clinically grounded responses from curated case scripts under explicit information-disclosure constraints to preserve realism. All interactions are then logged and evaluated by an Evaluation Agent, which compares the student’s inquiries with predefined checklists and applies an LLM-based judge to assess completeness and reasoning accuracy. At the end of each consultation, the Evaluation Agent produces structured feedback summarizing strengths and areas for improvement.
3.2 SPBench: Benchmark for Virtual SP
To enable systematic and reproducible evaluation of LLM-based SP, we constructed SPBench, a benchmark dataset derived from authentic SP–physician dialogues.
@@ 3.2 SPBench: Benchmark for Virtual SP
Case Source and Scope
We collected 58 real clinical cases. From the doctor-patient dialogue records associated with these cases, we processed and extracted a dataset containing 3,208 question-and-answer pairs. The source materials include the Manual for Writing Standardized Patient Cases [^1] https://weread.qq.com/web/bookDetail/1a432e10811e38145g0194ab and the A Practical Tutorial for Standardized Patients [^2] https://www.pumpedu.com/home-shop/7125.html. All records were fully anodized and used solely for academic research. They were also reviewed by experienced clinicians (see Appendix 9 for expert qualifications) to ensure accuracy and representativeness. As shown in Figure 2 and Figure 3, the specialty distribution of SPBench spans 14 clinical departments. General Surgery is the most represented specialty, accounting for 13.79% of the cases, followed by other departments such as Cardiology, Respiratory Medicine, and Psychiatry. The age of the patients ranges from children to the elderly, with the majority concentrated in the 21-60 age group. Regarding gender, the dataset is composed of 35 female cases, representing 60.3% of the total, and 23 male cases, representing the remaining 39.7%.
Data Structure
SPBench provides two core types of content: 1) Case Scripts: These provide standardized patient histories and physical examination findings for each case; and 2) Authentic Dialogue Records: These are turn-by-turn transcripts of conversations between physicians and human SP. The records serve as a benchmark to compare the conversational quality of an LLM-based virtual SP against a human SP for the same case.
Quality Control
To ensure accuracy and reliability, all dialogue transcripts were digitized from the two books in Section 3.2 using Optical Character Recognition (OCR) and then manually proofread by three fifth-year medical students. They corrected OCR-induced errors such as typos and punctuation to ensure fidelity to the original source material. This process preserves the authenticity and consistency required for benchmarking.
@@ 3.3 Evaluation Protocol
To systematically assess the performance of LLM as Standardized Patients, we (i) define what to measure via eight core evaluation criteria; (ii) validate the reliability of our automated scoring against clinical experts; (iii) conduct fair comparisons between human SP, EasyMED, and an external framework under identical inputs and scoring; and (iv) perform component-level ablations within EasyMED to analyze where gains come from. The following subsections detail each component.
Evaluation Metric
In collaboration with three clinical experts (see Appendix 9 for qualifications), we defined eight core dimensions for evaluating the virtual patient’s response quality, as detailed in Table 3. During the evaluation, experts rated performance for each dimension on a 5-point scale. These scores were then linearly transformed to a 100-point scale for reporting to enhance the granularity and readability of the results.
Expert Rating and Correlation Analysis
Considering the high cost of large-scale manual annotation, we first used a well-designed evaluation prompt (Appendix I.4) and employed GPT-4o as an automated evaluator to score all results. To verify its reliability (see Appendix 9 for qualifications), we randomly selected 30 samples for blind review by two clinical experts who were unaware of the GPT-4o ratings. The automated scores correlated strongly and significantly with the experts’ average ratings (Pearson’s $r=0.79$; Table 2). Although GPT-4o mean score (78.5) was consistently stricter than the experts’ mean (83.9), this high alignment supports its use as a reliable proxy for large-scale evaluation.
Automated Professional A Professional B
Average Score 78.5 86.3 81.5
Standard Deviation 8.8 9.3 9.8
Correlation 0.79
Table 2: Inter-Rater Reliability between an Automated Evaluator and Human Experts
Evaluation Criterion Description
Query Comprehension Assesses the ability to accurately understand the physician’s questions, grasp the core intent, and provide relevant answers without misinterpretation.
Case Consistency Evaluates whether responses are factually consistent with the predefined patient case, including symptoms, medical history, and key facts. Checks for contradictions or information not present in the case.
Controlled Disclosure Measures the ability to provide information passively and appropriately, answering only what is asked without volunteering unsolicited details. Focuses on conciseness and relevance.
Response Completeness Assesses whether responses are comprehensive and address all parts of a physician’s query, avoiding the omission of crucial information from the case profile.
Logical Coherence Evaluates the internal logical consistency of the information provided. Ensures that reported symptoms and their characteristics (e.g., location, duration) are coherent from a patient’s perspective.
Language Naturalness Assesses whether the communication style is natural and patient-like, using plain language and avoiding overly technical medical jargon. Focuses on the authenticity of the linguistic style.
Conversational Consistency Measures the ability to maintain consistent information across a multi-turn dialogue. Checks for self-contradictions and ensures stability throughout the interaction.
Patient Demeanor Evaluates the ability to maintain a patient-like demeanor, demonstrating appropriate patience, cooperation, and emotional tone, even when faced with repetitive or clarifying questions.
Table 3: Definitions of the eight evaluation criteria for assessing the quality of the conversational agent.
Comparison between human and LLM-based SP
To compare the responses of LLM-based SP with those of a human SP under identical context, we first isolated the physician’s questions from the dialogue records in SPBench. We then used these exact questions and patient case as inputs for the LLM-based SP. This method allowed us to directly compare the responses generated by our virtual patients against those of the human SP under identical conditions.
Comparison with External Frameworks
Beyond comparing against human SP under identical prompts, we further benchmarked our system against an open-source LLM-based SP framework, EvoPatient du2024llms. Using the same SPBench physician questions and case profiles, and the same automated evaluator and scoring protocol, basic settings are summarized in Appendix F. we report the system-level scores in Table 4.
System Configuration Score
Human SP 97.33
EasyMED (ours) More Agent 96.98
EvoPatient Authors setup 93.33
Table 4: Comparison of different frameworks on SPBench under the same evaluation protocol.
Ablation on EasyMED Components
We conducted an ablation study to quantify the contribution of key components in our virtual SP design. For each model, we started from a baseline configuration and incrementally added components: (1) Baseline no chain-of-thought or intent recognition; (2) CoT adds explicit step-by-step reasoning to the baseline; (3) Intent Recognition further adds an intent-recognition module on top of CoT.
Model Name Baseline COT Intent Recognition
Qwen3-8B 81.12 81.95 81.64
Qwen3-32B 88.87 91.89 94.08
DeepSeek-R1 93.11 94.90 95.72
GPT-4o 89.54 90.08 93.70
Gemini 2.5 Pro 95.04 95.77 96.98
Table 5: Effect of CoT and intent recognition modules across different models.
Table 4 shows that EasyMED matches human SP performance (96.98 vs. 97.33) and clearly outperforms the external framework EvoPatient (93.33). These results indicate that our multi-agent design and evaluation-aligned prompting produce higher-fidelity SP behaviors under identical inputs and scoring.
As shown in Table 5, adding COT to the baseline improves most models, and layering intent recognition on top yields further gains for medium-to-large models (e.g., Qwen3-32B, DeepSeek-R1, GPT-4o, Gemini 2.5 Pro). However, these advanced components are capacity-sensitive: for a small model such as Qwen3-8B, intent-guided CoT can introduce over-constraints and slightly reduce the score relative to plain CoT.

### v2

# Human or LLM as Standardized Patients? A Comparative Study in Medical Education

@@ Abstract
Standardized patients (SPs) are indispensable for clinical skills training but remain expensive and difficult to scale. Although large language model (LLM)–based virtual standardized patients (VSPs) have been proposed as an alternative, their behavior remains unstable and lacks rigorous comparison with human standardized patients. We propose EasyMED, a multi-agent VSP framework that separates case-grounded information disclosure from response generation to support stable, inquiry-conditioned patient behavior. We also introduce SPBench, a human-grounded benchmark with eight expert-defined criteria for interaction-level evaluation. Experiments show that EasyMED more closely matches human SP behavior than existing VSPs, particularly in case consistency and controlled disclosure. A four-week controlled study further demonstrates learning outcomes comparable to human SP training, with stronger early gains for novice learners and improved flexibility, psychological safety, and cost efficiency.
@@
A Comparative Study in Medical Education
@@ 1 Introduction
Clinical reasoning and doctor–patient communication are essential skills in medical education Cleland and Durning (2022). Their development relies on repeated, interactive practice in realistic clinical settings. Standardized patients, trained actors who consistently portray predefined clinical cases, are widely regarded as the gold standard for teaching and assessing these skills, particularly in Objective Structured Clinical Examinations (OSCEs) Sayers et al. (2024); Ma et al. (2023). While SP-based training enables safe and authentic clinical encounters, human SP programs are costly, labor-intensive, and difficult to scale, which limits training frequency and accessibility Zendejas et al. (2013). Consequently, large language model (LLM) based virtual standardized patients (VSP) have emerged as a promising scalable alternative Du et al. (2024); Ye and Tang (2025), due to their strong dialogue capabilities and broad world knowledge.
Despite recent progress, it remains unclear whether LLM-based VSP can support clinical skills training at a level comparable to human standardized patients. This question is difficult to answer due to persistent gaps in system design and evaluation that are misaligned with real SP training practice. Most VSP frameworks conflate inquiry interpretation with response generation, leading to premature information disclosure, cross-turn instability, and limited support for intent-aware instructional feedback Du et al. (2024); Ye and Tang (2025); Sirdeshmukh et al. (2025). Existing evaluations are largely coarse-grained, relying on synthetic dialogues or outcome-level metrics rather than authentic human SP–doctor interactions Fan et al. (2023); Waisberg et al. (2024). Moreover, systematic long-term comparisons with human standardized patients under matched training conditions remain rare Liu et al. (2025); Bodonhelyi et al. (2025), leaving the educational effectiveness of LLM-based virtual patients insufficiently validated.
SP Patient Fidelity Interaction Coherence Pedagogical Awareness Real-world User Study
Human SP -
SimPatient Steenstra et al. (2025)
EvoPatient Du et al. (2024)
CureFun Li et al. (2024b)
EasyMED (Ours)
Patient Agent Auxiliary Agent Evaluation Agent (Sec. 6)
intent-conditioned disclosure factorized patient simulation trajectory-level feedback
Table 1: Comparison of various standardized patient systems across three desiderata defined in Section 2, based on what is explicitly reported in the original papers. SimPatient evaluates realism primarily via qualitative user studies and adopts end-to-end patient response generation, while providing utterance-level reflective feedback. EvoPatient emphasizes emergent realism through agent co-evolution without explicit control over information disclosure or learner-facing instructional feedback. CureFun focuses on educational usefulness without controlled comparison to human standardized patients and performs end-to-end patient simulation, while offering post-session learning summaries. EasyMed embeds patient fidelity, interaction coherence, and pedagogical awareness directly into its architecture via intent-conditioned disclosure calibrated by SPBench (Sec. 5), factorized patient simulation using Auxiliary Agent (Sec. 3.3), and trajectory-level evaluation in Evaluation Agent (Sec. 3.4).
Multi-agent VSP To address the limitations identified above, We propose EasyMED, a controllable multi-agent framework that models virtual SP training as a structured, interactive process. EasyMED decouples patient simulation, intent recognition, and evaluation into coordinated agents, enabling intent-conditioned information disclosure, stable cross-turn behavior, and checklist-based instructional feedback. This design directly supports patient fidelity, interaction coherence, and pedagogical awareness in virtual SP training.
SP Benchmark To support reproducible and interaction-level evaluation, we further introduce SPBench, a benchmark constructed from authentic standardized patient-doctor dialogues spanning 14 medical specialties and eight expert-defined evaluation criteria Fan et al. (2023); Sirdeshmukh et al. (2025). Unlike existing benchmarks that rely on synthetic dialogues or outcome-level scores, SPBench uses human SP interaction trajectories as reference to quantitatively compare virtual and human standardized patient behavior. Using SPBench, we compare EasyMED and representative existing VSPs against human SP interaction trajectories, and find closer alignment with human SP behavior, particularly in controlled disclosure and cross-turn consistency.
Real-world User Study To assess educational effectiveness in real training settings, we conduct a four-week controlled study comparing EasyMED with human standardized patient training under matched content and scoring rubrics, directly addressing whether LLM-based virtual patients can achieve training effectiveness comparable to human SPs.
Our contributions are threefold: (1) we propose EasyMED, a multi-agent virtual standardized patient framework that enables controllable, interpretable patient simulation aligned with clinical training workflows; (2) we propose SPBench, a human-grounded benchmark that enables quantitative, interaction-level comparison between virtual and human standardized patient behavior; (3) we present a controlled, longitudinal user study comparing LLM-based and human standardized patient training under matched conditions.
Figure 1: Overview of the multi-agent architecture of the virtual standardized patient system, consisting of a Patient Agent, an Intent Recognition Agent, and an Evaluation Agent.
2 Background
VSPs are increasingly adopted for clinical training due to their scalability and accessibility. However, their educational value hinges not only on their ability to simulate patient dialogue, but also on whether they can faithfully reproduce human standardized patient behavior (see Desideratum I), remain stable and controllable across extended interactions (see Desideratum II), and actively support clinical learning feedback (see Desideratum III). In practice, current VSPs often fall short in one or more of these aspects, limiting their reliability as training tools, see Table 1.
Desideratum I.
Patient Fidelity. VSPs should exhibit behaviors and response patterns comparable to those of human standardized patients.
Achieving such fidelity requires evaluation protocols that enable direct experimental comparison with human SPs, rather than relying solely on subjective user satisfaction. However, most prior studies on VSPs primarily adopt qualitative evaluations based on structured interviews and satisfaction surveys Steenstra et al. (2025); Du et al. (2024); Li et al. (2024b). As a result, direct comparisons with human SPs remain rare, and the educational effectiveness of LLM-based VSPs relative to traditional human SP training is still unclear Simzine (2025).
Desideratum II.
Interaction Coherence. VSPs must maintain consistent clinical states and controlled information disclosure across multi-turn interactions.
Interaction coherence requires separating what clinical information is revealed from how it is expressed, so as to prevent case drift and unintended information leakage across turns. However, most existing VSPs generate responses in an end-to-end manner without explicitly modeling this separation Steenstra et al. (2025); Li et al. (2024b). As a result, although responses may appear locally coherent, longer interactions often exhibit cross-turn instability and uncontrolled disclosure, undermining the reliability of VSPs for medical education.
Desideratum III.
Pedagogical Awareness. VSP-based training systems should be aware of the learner’s educational objectives and interaction process, enabling fine-grained instructional feedback that supports clinical learning rather than merely simulating patient behavior.
Pedagogical awareness requires VSPs to go beyond passive patient simulation and actively support learning by monitoring the learner’s inquiry process and identifying opportunities for guidance. Grounding feedback in the interaction trajectory allows such systems to highlight missing, redundant, or inappropriate clinical questions and support reflective learning. However, most existing VSPs function primarily as conversational agents and lack explicit representations of clinical intent or inquiry coverage, limiting their ability to provide meaningful instructional feedback Steenstra et al. (2025); Du et al. (2024).
3 EasyMED: A Multi-Agent VSP Framework
Criterion Abbr. Description
Query Comprehension QC Accurate understanding of the physician’s question and its intent without misinterpretation
Case Consistency $C$ Faithfulness to the predefined patient case, without contradictions or unsupported facts.
Controlled Disclosure CD Providing only requested information, avoiding unsolicited or premature disclosure.
Response Completeness RC Fully addressing all aspects of the physician’s query without omitting essential case information.
Logical Coherence LC Internal logical consistency of responses, ensuring symptoms and attributes remain coherent.
Language Naturalness LN Use of natural, patient-like language while avoiding unnecessary medical jargon.
Conversational Consistency CS Consistency of information across dialogue turns, avoiding self-contradictions.
Patient Demeanor PD Maintaining an appropriate patient-like emotional tone, including cooperation and stability.
Table 2: Definitions of the eight evaluation criteria used in SPBench.
3.1 Workflow of EasyMED
Philosophy of EasyMED
Section 2 identifies three desiderata for virtual standardized patients—patient fidelity, interaction coherence, and pedagogical awareness—that are difficult to achieve with end-to-end simulators. EasyMED treats these desiderata as explicit design constraints and maps them to concrete architectural choices: intent-conditioned response generation to preserve patient fidelity, decoupled case-grounded information access and surface realization to ensure interaction coherence, and trajectory-level retention for checklist-based educational feedback. This principled mapping naturally motivates a factorized, multi-agent design.
EasyMED implements a factorized workflow with two phases: consultation and evaluation, as shown in Figure 1. Phase 1 is realized by the Auxiliary and Patient Agents for intent recognition and patient simulation, while Phase 2 is conducted by the Evaluation Agent for trajectory-level assessment and feedback.
Phase 1: Consultation During consultation, the interaction unfolds as a multi-turn dialogue
$\mathcal{D}=\{(q_{1},r_{1}),\ldots,(q_{T},r_{T})\},$ (1)
where $q_{t}$ is the learner’s question and $r_{t}$ the patient response at turn $t$. At each turn, EasyMED first infers a standardized clinical intent
$i_{t}=A(q_{t},H_{t-1}),$ (2)
and then generates a case-grounded response
$r_{t}=P(i_{t},q_{t},H_{t-1}\mid E).$ (3)
By factorizing intent recognition and response generation, EasyMED enables inquiry-conditioned disclosure and stable multi-turn patient behavior.
Phase 2: Evaluation After the consultation ends, the Evaluation Agent reviews the full dialogue trajectory $\mathcal{D}$ and compares the recognized intents and elicited facts against expert-defined case checklists to produce structured feedback.
3.2 Auxiliary Agent
The Auxiliary Agent addresses a core limitation of existing virtual standardized patients by explicitly modeling the learner’s clinical inquiry rather than relying on end-to-end text generation. It maps each learner question to a predefined clinical intent (e.g., Chief Complaint or Onset), abstracting away surface-level linguistic variation. This standardized intent representation serves as a control signal for downstream patient simulation, ensuring that responses are conditioned on inquiry type rather than phrasing, thereby enabling controlled information disclosure and stable patient behavior across multi-turn interactions.
3.3 Patient Agent
The Patient Agent simulates patient behavior while enforcing case fidelity and disclosure constraints. Given an inferred clinical intent, it first retrieves the corresponding fact from a structured electronic health record that defines the patient’s ground-truth case information, and then generates a natural language response conditioned on both the retrieved fact and a predefined patient persona (e.g., anxiety or hesitation). This separation between fact selection and surface realization enables controlled information disclosure while preserving natural conversational flow.
3.4 Evaluation Agent
The Evaluation Agent acts as a post-hoc pedagogical observer. Rather than intervening during conversation, it reviews the interaction history once the session ends. By comparing recognized intents and elicited facts against the standard case checklist, the agent generates structured feedback to highlight gaps like Missed Inquiries. This provides students with actionable feedback, addressing the evaluation limitations discussed in Section 2.
4 SPBench: Benchmark for Virtual SP
4.1 The Philosophy of SPBench
Existing medical benchmarks primarily assess static knowledge or aggregate outcomes and fail to capture interactive patient behavior (e.g., MMLU Hendrycks et al. (2021), MedQA Jin et al. (2021)), forcing VSP evaluation to rely on synthetic or interview-based data Fan et al. (2023); Waisberg et al. (2024). SPBench fills this gap by grounding evaluation in authentic human SP–doctor dialogue trajectories and adopting a standardized protocol with eight clinically motivated dimensions that assess both turn-level response quality and session-level behavioral consistency.
4.2 Data Curation
Data Structure SPBench contains two main parts: (1) Patient Profile: These describe the patient’s main complaint, symptoms, history, and background; and (2) Authentic Dialogue Records: turn-level transcripts showing how trained human SP respond to clinical questions in real instructional settings. We use these records as a standard. They allow us to compare the AI’s performance against a human actor handling the same case.
System QC $C$ CD RC LC LN CS PD Overall
LLMs
Qwen3-8B 77.76 78.64 84.26 85.74 83.37 72.14 85.74 81.31 81.12
Qwen3-32B 88.37 88.37 89.31 89.31 88.99 85.89 89.91 89.61 88.87
DeepSeek-R1 91.06 93.08 93.42 97.10 96.10 85.38 93.69 95.05 93.11
GPT-4o 75.87 89.24 88.89 98.11 91.31 91.66 90.62 90.62 89.54
Gemini 2.5 Pro 94.72 95.05 96.04 95.69 94.39 93.21 96.04 95.27 95.04
Prompt Strategy + Agent Framework
Gemini 2.5 Pro + CoT 96.61 94.98 96.63 95.33 93.99 95.59 97.61 95.61 95.77
Human SP (reference) 95.71 96.47 98.53 97.85 98.18 95.10 98.56 98.22 97.33
EvoPatient 91.87 96.29 94.59 95.77 92.67 90.49 92.32 92.74 93.33
EasyMED (ours) 97.17 97.23 98.18 97.11 95.03 97.48 97.98 95.73 96.98
Table 3: Overall and per-dimension performance evaluation on SPBench. Human SP serves as the gold standard reference. The best-performing LLM-based system is highlighted in bold.
Case Source and Scope SPBench is derived from two widely used training books for Standardized Patients—the Manual for Writing Standardized Patient Cases [^1] https://www.pmph.com/ and the A Practical Tutorial for Standardized Patients [^2] https://www.pumpedu.com/home-shop/7125.html. From these resources, we collected 3,208 question–answer pairs used in real doctor–patient interactions. We cleaned and refined this data into 58 separate patient cases. Each case includes a profile and its dialogue records. The dataset is intended exclusively for academic research and evaluation purposes.
Two clinical experts (Appendix B) checked each case to ensure anonymity, realism, and pedagogical usefulness. As shown in Figures 3 and 2, SPBench covers 14 medical fields.
Quality Control To ensure accuracy and reliability, we scanned the books using Optical Character Recognition (OCR). Then, three senior medical students manually verified the extracted text and corrected typographical and punctuation errors. This step ensures the data faithfully matches the original books, which is necessary for a reliable benchmark.
@@ 4.3 Evaluation Protocol
Evaluation Metric SPBench evaluates VSP by assessing both turn-level response quality and session-level interactional behavior. Unlike existing benchmarks that rely on static knowledge tests or single aggregate scores, our evaluation decomposes patient performance into multiple clinically interpretable dimensions.
Specifically, we define eight evaluation criteria (Table 2) in collaboration with three clinical experts (see Appendix B). These criteria capture complementary aspects of patient simulation, including accurate understanding of clinical inquiries, adherence to the predefined case, controlled disclosure of information, and consistency across dialogue turns. Each criterion is independently rated on a 5-point Likert scale, and scores are linearly rescaled to a 100-point scale for reporting.
Input Standardization To ensure fair and reproducible evaluation, SPBench uses authentic questions extracted from real-world doctor–patient dialogues. For each clinical case, we isolate the sequence of questions asked by the human physician and present this exact sequence to each model. This design eliminates variability introduced by different prompting styles and ensures that all virtual standardized patients are evaluated under identical input conditions.
5 Evaluation on SPBench
This section analyzes divergences between virtual and human standardized patients in multi-turn interactions and assesses how EasyMED reduces these gaps relative to human SP behavior.
Overall Performance  Table 3 summarizes overall and per-dimension performance on SPBench. Human standardized patients achieve the highest reference score (97.33), reflecting stable case portrayal and appropriate information disclosure across turns. EasyMED closely matches human performance (96.98), with strongest gains on interaction-critical dimensions (CC, CD, CS, PD) that directly align with standardized patient requirements. By contrast, although several frontier LLMs perform well on LN and RC, they show larger variance on CC and CD, indicating unstable case grounding and inconsistent inquiry-conditioned disclosure.
Sensitivity on Prompting Strategies  To disentangle the sources of performance differences, we first examine LLM baselines under a unified prompting scheme. Although large models such as Gemini 2.5 Pro achieve relatively balanced scores, consistent weaknesses remain in controlled disclosure and cross-turn stability. We further evaluate common prompting strategies using a fixed backbone (Gemini 2.5 Pro). CoT prompting improves reasoning transparency and modestly increases CC and RC. However, it also amplifies verbosity and unsolicited explanation, leading to reduced CD scores. Overall, prompting mainly influences response articulation rather than information control, and is insufficient to enforce standardized patient behavior in multi-turn interactions.
Ablation Study on Auxiliary Agent  We next examine the effect of the Auxiliary Agent in EasyMED, which decouples intent recognition from response generation. As shown in Table 3, this component yields the largest gains on CC, CD, CS, and PD. By mapping learner queries to standardized clinical intents, the Auxiliary Agent provides an explicit control signal that constrains information access, mirroring how human standardized patients condition responses on inquiry type rather than full case narratives. These gains exceed those from prompt engineering alone, highlighting the dominant role of architectural control in stabilizing multi-turn patient behavior.
6 Real-world Evaluation in Medical Education
While SPBench evaluates interaction-level fidelity, it does not directly capture educational effectiveness. We therefore assess EasyMED in a real training setting through a controlled user study, comparing it with human standardized patient training in terms of learning outcomes, learner experience, and practical feasibility.
Study Period Timeline Group A Group B
Baseline Week 0 Pre-Test
Period 1 Weeks 1-2 EasyMED Training Human SP Training
End of Week 2 Mid-Test
Period 2 Weeks 3-4 Human SP Training EasyMED Training
End of Week 4 Final Test & Questionnaire
Table 4: Experimental design of the four-week study, showing the sequence of training interventions and assessments for Group A and Group B.
6.1 Experimental Design
We adopted a randomized crossover design to enable within-subject comparison while controlling for baseline differences. Each participant experienced both EasyMED and human SP training in different phases, which allows analysis of overall learning gains as well as phase-specific effects attributable to each modality (Table 4).